# Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning

### v1

@@ Abstract
Few-shot named entity recognition can identify new types of named entities based on a few labeled examples. Previous methods employing token-level or span-level metric learning suffer from the computational burden and a large number of negative sample spans. In this paper, we propose the Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning (MsFNER), which splits the general NER into two stages: entity-span detection and entity classification. There are 3 processes for introducing MsFNER: training, finetuning, and inference. In the training process, we train and get the best entity-span detection model and the entity classification model separately on the source domain using meta-learning, where we create a contrastive learning module to enhance entity representations for entity classification. During finetuning, we finetune the both models on the support dataset of target domain. In the inference process, for the unlabeled data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN. We conduct experiments on the open FewNERD dataset and the results demonstrate the advance of MsFNER.
Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning
Peipei Liu1,2, Gaosheng Wang1,2, Ying Tong3, Jian Liang3, Zhenquan Ding1thanks: The corresponding author, Hongsong Zhu1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Jiangsu Provincial Department of Public Security, China {liupeipei, wanggaosheng, dingzhenquan, zhuhongsong}@iie.ac.cn
1 Introduction
@@
Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations Yadav and Bethard (2018); Li et al. (2022). Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition (NER) task with sufficient training data Lample et al. (2016); Ma and Hovy (2016), they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes. In this case, researchers have proposed the few-shot named entity recognition (Few-shot NER) to explore entity recognition with limited labeled data Fritzler et al. (2019). Few-shot NER enables existing models to quickly transfer learned knowledge and adapt to new domains or entity classes.
@@
Specifically, the Few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ Snell et al. (2017). $S_{target}$ represents the support dataset comprising $N$ entity types ($N$-way), and each type is exemplified by $K$ annotated examples ($K$-shot). The $Q_{target}$ is the query dataset with the same entity types as the support dataset. Few-shot paradigm can offer a flexible and cost-effective solution to the adaptability challenge, making it a focal point of research to enhance the performance of NER systems in scenarios with limited labeled data or emerging entity types.
@@
Currently, there are two mainstream researches for Few-shot NER: token-level Fritzler et al. (2019); Hou et al. (2020); Yang and Katiyar (2020); Das et al. (2022) and span-level metric learning methods Wang et al. (2022a); Yu et al. (2021); Ma et al. (2022). In token-level methods, each token is assigned an entity label based on its distances from the prototypes of entity classes or the support tokens. However, these approaches often have high computational costs and fail to maintain the semantic integrity of entity tokens, leading to increased susceptibility to interference from non-entity markers. On the other hand, although span-level methods can mitigate the partial issues associated with token-level approaches by evaluating entities as spans, all possible spans are enumerated would result in an N-square complexity and an increase in noise from a large number of negative samples.
@@
In this paper, we propose a hybrid multi-stage decoding approach for few-shot NER with contrastive learning (i.e., MsFNER). Specifically, MsFNER splits the general NER into two stages and this paper describes them by three processes: training process, finetuning process, and inference process. The training process consists of two objectives: training the best entity span detection model and the best entity classification model separately based on the supervised support dataset of source domain. In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)Finn et al. (2017), and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNetMa et al. (2022). In the finetuning process, we finetune the trained entity detection model and entity classification model on the support dataset of target domain with the same pattern as training process. In the inference process, we deploy the KNN and two finetuned models on the query dataset of target domain to predict entity spans and their types, and there are four phases for prediction. In the first phase, we build a key-value datastore $\mathcal{D}_{knn}$ on the support dataset $S_{target}$, where key is the entity representation and value is the corresponding label. Subsequently, we apply the entity detection model to get entity spans. With the entity detection model, we could reduce the impact of negative samples and reduce computational complexity. After that, we pass the representations of detected entity spans into the entity classification model to produce their predicted type distribution $p_{soft}$ while we also feed the representations into KNN to obtain the predicted results $p_{knn}$ based on the $\mathcal{D}_{knn}$. In the final phase, we combine $p_{soft}$ and $p_{knn}$ to obtain final types for prediction entities.
@@
  • Experimental results show that our model achieves new SOTA performance compared with previous works. We also evaluate the few-shot NER task on LLM ChatGPT and the experimental results show that our model outperforms ChatGPT in terms of performance and efficiency.
2 Methodology
@@
2.1 Training Process
@@
Figure 1: The schematic diagram of the training process.
2.1.1 Entity Span Detection (ESD)
@@
For a given sentence $x$ with $n$ tokens $x$=$(x_{1},x_{2},...,x_{n})$ in $D_{source}$, we first encode it using the pre-trained language model BERT. After that, we can obtain the contextualized representations $h=(h_{1},h_{2},...,h_{n})$ for all tokens.
@@
where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions.
@@
2.1.2 Entity Classification
@@
[…]
@@
where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot ($K$>1) but the squared euclidean distance for $K$-shot ($K$=1) referring to Das et al. (2022), and $\tau$ is a temperature hyperparameter.
After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning Chen et al. (2020); Khosla et al. (2020); Liu et al. (2023). As a consequence, we use the entity representation $\hat{e}_{k}$ to compute the prototype $c_{t}(\mathcal{S})$ of $t_{th}$ type Snell et al. (2017), and predict the type distribution of $e_{k}$ by:
$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii\in\phi}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)}$ (5)
@@
2.2 Finetuning Process
Model finetuning aims to enable the model to adapt to new knowledge types during the domain transfer process. We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same method as training process.
2.3 Inference Process
@@
(2) We input the query sentence $x^{\prime}$ into the finetuned best model for entity span detection (ESD) to get all entities. Specifically, we apply the Viterbi algorithm Forney (1973) for decoding, and derive the entities from the maximum sequence probability of Eq.2:
@@
Table 1: F1 scores with standard deviations on FewNERD. †denotes the results reported in Ma et al. (2022). ‡are the results we produce by LLM. * means the abbreviation of ‘contrastive learning’. The best results are in bold.
(3) Firstly, we compute the representation $\hat{e}^{\prime}_{k}$ of each detected entity ${e}^{\prime}_{k}$ by Eq.3. Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ Wang et al. (2022b), and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.5.
@@
where the $\lambda$ is a hyper-parameter that makes a balance between KNN and softmax.
3 Experiments
3.1 Experiments Setup
The details can be found at A.1.
3.2 Main results
Table 1 illustrates the main results of different methods on FewNERD dataset. The results demonstrate that MsFNER significantly outperforms all the previous state-of-the-art approaches. MsFNER attains average F1 improvements of 2.65 and 4.44 on INTRA and INTER, respectively, compared to the previous best method, MAML-ProtoNet. In the 5-way 5-10 shot setting on the INTER dataset, we achieve the most substantial improvement over the MAML-ProtoNet method, reaching 7.79 points. Additionally, we perform the few-shot task employing ChatGPT, where MsFNER surpasses ChatGPT in F1 performance by an average score of 8.465 in the 5-10 shot setting. Furthermore, MsFNER surpasses ChatGPT by an average of 10.13 percentage points in the 1-2 shot setting in the INTER dataset. Overall, regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance.
3.3 Ablation Study
To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time to observe the performance of our model. The results are shown at the bottom of Table 1. Firstly, we remove the contrastive learning module. In the 1 shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset. We can see that the removal of contrastive learning results in a decrease of 0.905 average score, which indicates the significant impact and necessity of contrastive learning for enhancing representation. Secondly, we remove the KNN module and reserve the contrastive learning module. The removal of KNN drops 0.523 average score. We can see that the impact of contrastive learning and KNN increases with the number of $N$-way and $K$-shot. Removing either of them will lead to a significant decrease in the performance of MsFNER.
3.4 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. More details can be found at A.3.
4 Conclusion
In this paper, we propose the MsFNER, aiming to improve the performance and efficiency of few-shot NER. The MsFNER can be achieved by three stages: training, finetuning and inferring. We evaluate MsFNER on the open FewNERD dataset, and results show the advance of MsFNER compared with previous few-shot NER methods and LLM.
5 Limitations
  • In practical, the effectiveness of Few-shot NER is seriously affected due to the data sparsity and data imbalance.
  • The abundant external knowledge is very helpful for few-shot NER, which is worth exploring and studying.
6 Ethical Considerations
All the data we get is the open source, and there is no theft. We guarantee the originality, research ethics, societal impact and reproducibility of our work.
Appendix A Appendix
A.1 Experiments Setup

### v2

@@ Abstract
Few-shot named entity recognition can identify new types of named entities based on a few labeled examples. Previous methods employing token-level or span-level metric learning suffer from the computational burden and a large number of negative sample spans. In this paper, we propose the Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning (MsFNER), which splits the general NER into two stages: entity-span detection and entity classification. There are 3 processes for introducing MsFNER: training, finetuning, and inference. In the training process, we train and get the best entity-span detection model and the entity classification model separately on the source domain using meta-learning, where we create a contrastive learning module to enhance entity representations for entity classification. During finetuning, we finetune the both models on the support dataset of target domain. In the inference process, for the unlabeled query data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN. We conduct experiments on the open FewNERD dataset and FewAPTER dataset, the results demonstrate the advance of MsFNER.
I Introduction
@@
Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations [1, 2]. Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition task with sufficient training data [3, 4], they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes. In this case, researchers have proposed the few-shot named entity recognition (few-shot NER) to explore entity recognition with limited labeled data [5]. Few-shot NER enables existing models to quickly transfer learned knowledge and adapt to new domains or entity classes.
@@
Specifically, the few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ [6]. $S_{target}$ represents the support dataset comprising $N$ entity types ($N$-way), and each type is exemplified by $K$ annotated examples ($K$-shot). The $Q_{target}$ is the query dataset with the same entity types as the support dataset. Few-shot paradigm can offer a flexible and cost-effective solution to the adaptability challenge, making it a focal point of research to enhance the performance of NER systems in scenarios with limited labeled data or emerging entity types.
@@
Currently, there are two mainstream researches for few-shot NER: token-level [5, 7, 8, 9] and span-level metric learning methods [10, 11, 12]. In token-level methods, each token is assigned an entity label based on its distances from the prototypes of entity classes or the support tokens. However, these approaches often have high computational costs and fail to maintain the semantic integrity of entity tokens, leading to increased susceptibility to interference from non-entity markers. On the other hand, although span-level methods can mitigate the partial issues associated with token-level approaches by evaluating entities as spans, all possible spans are enumerated would result in an N-square complexity and an increase in noise from a large number of negative samples.
@@
In this paper, we propose a hybrid multi-stage decoding approach for few-shot NER with contrastive learning (i.e., MsFNER). Specifically, MsFNER splits the general NER into two stages and this paper describes them by three processes: training process, finetuning process, and inference process. The training process consists of two objectives: training the best entity span detection model and the best entity classification model separately based on the supervised support dataset of source domain. In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)[13], and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNet[12]. In the finetuning process, we finetune the trained entity detection model and entity classification model on the support dataset of target domain with the same pattern as training process. In the inference process, we deploy the KNN and two finetuned models on the query dataset of target domain to predict entity spans and their types, and there are four phases for prediction. In the first phase, we build a key-value datastore $\mathcal{D}_{knn}$ on the support dataset $S_{target}$, where key is the entity representation and value is the corresponding label. Subsequently, we apply the entity detection model to get entity spans. With the entity detection model, we could reduce the impact of negative samples and reduce computational complexity. After that, we pass the representations of detected entity spans into the entity classification model to produce their predicted type distribution $p_{soft}$ while we also feed the representations into KNN to obtain the predicted results $p_{knn}$ based on the $\mathcal{D}_{knn}$. In the final phase, we combine $p_{soft}$ and $p_{knn}$ for predicting entity types.
@@
  • Experimental results show that our model achieves new SOTA performance compared with previous works. We also evaluate the few-shot NER task on LLM ChatGPT, and MsFNER can have the comparable comprehensive results with ChatGPT in terms of performance and efficiency.
II Methodology
@@
Figure 1: The schematic diagram of the training process.
II-A Training Process
@@
Entity Span Detection (ESD)
II-A1 ESD Method
@@
[…]
@@
where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions, $y^{\prime}$ is the possible label sequence.
II-A2 Training
@@
Specifically, the model first randomly samples a batch of task data $\mathcal{T}=\{(S_{source}^{(i)},Q_{source}^{(i)})\}_{i=1}^{bz}$ from the source domain data $D_{source}$, where the batch size is $bz$. Next, the model is trained separately on $\mathcal{T}_{s}=\{S_{source}^{(i)}\}_{i=1}^{bz}$ and $\mathcal{T}_{q}=\{Q_{source}^{(i)}\}_{i=1}^{bz}$, which we refer to as inner-loop update (Task-level adaptation) and outer-loop update (Meta-optimization), respectively. An overview of the training process is presented in Algorithm 1.
In the inner-loop update, for each learning episode in meta-learning, the model performs one gradient update on the task data $S_{source}^{(i)}$ to update the parameters:
$\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$ (3)
where $\theta=\{\theta^{encoder},\theta^{crf}\}$ represents the model’s learnable parameters, $\theta_{i}^{\prime}$ is the updated parameter after one gradient update on the task data $S_{source}^{(i)}$, and $\alpha$ is the learning rate used to minimize the loss function (Equation (1)).
After the inner-loop update, the model will fully utilize the labeled query set $\mathcal{T}_{q}$ from the source domain for the outer-loop update. This update differs from the inner-loop update’s model optimization approach (i.e., using the support set of one task in each learning episode for parameter optimization). Instead, it follows the conventional neural network training method, where the loss from all the training data in the batch is optimized together. The main purpose of using both inner and outer-loop update is to enhance the model’s generalization ability on new task data and prevent overfitting to the support set data. The optimization objective of the outer-loop update is:
$\min_{\theta}\sum\nolimits_{\mathcal{T}_{q}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ (4)
Although the loss calculation at this point involves $\theta_{i}^{\prime}$, the model’s goal remains to update and optimize $\theta$. The outer-loop update optimization is performed through Stochastic Gradient Descent (SGD), and therefore, the model parameter update process can be expressed as:
$\theta=\theta-\beta\sum\nolimits_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ (5)
where $\beta$ is the learning rate for SGD.
Algorithm 1 Training Process of Entity Span Detection Model Based on MAML
[…]
3:Randomly initialize $\theta$
4:while not done do
5:  Randomly select a batch of task data $\mathcal{T}$, where $\mathcal{T}\sim D_{source}$
6:  for all $\mathcal{T}$ do
7:   Select one task data $(S_{source}^{(i)},Q_{source}^{(i)})$ from $\mathcal{T}$
8:   Compute the gradient $\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$ based on the support set $S_{source}^{(i)}$
9:   Perform a gradient update: $\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$
10:   Compute the gradient $\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ based on the query set $Q_{source}^{(i)}$, and store the result
11:  end for
12:  Outer-loop update: $\theta=\theta-\beta\sum_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$
13:end while
Entity Classification (EC)
II-A1 EC Method
@@
In details, given a batch with N entities and the anchor index j $\in$ {1,2,…,N}, an entity-aware contrastive loss can be defined as follows:
@@
where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot (K>K>1) but the squared euclidean distance for $K$-shot (KK=1) referring to [9], and $\tau$ is a temperature hyperparameter.
After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning [14, 15, 16]. Subsequently, prototypes for different types of entities are constructed based on the enhanced entity features. Let $\mathcal{S}_{t}$ denote the set of entities belonging to the $t$-th type in the support dataset $S_{source}^{(i)}$ of a task during training. The prototype for the $t$-th type is obtained by calculating the average of all entity features in this type.
$c_{t}(\mathcal{S})=\frac{1}{\left|S_{t}\right|}\sum_{{e}_{m}\in S_{t}}\hat{e}_{m}$ (8)
As a consequence, we can predict the type distribution of $e_{k}$ by calculating the distance between its entity representation $\hat{e}_{k}$ and all type prototypes:
$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii=1}^{|\phi|}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)},t\in\{1,..,|\phi|\}$ (9)
where $\phi$ is the entity type set and $|\phi|$=NN, $d(\cdot,\cdot)$ is the distance function. So, the entity classification loss is formalized as:
$L_{EC}=-\sum_{k=1}^{NK}y_{k}log(p_{soft}({e}_{k}))$ (10)
where $y_{k}$ is the truth type of ${e}_{k}$, $L_{\mathit{EC}}$ is the loss which we optimize.
II-A2 Training
The training of the model in this stage (detailed in Algorithm 2) adopts the MAML-ProtoNet approach proposed by Ma et al. [12], which combines the MAML method with prototype network techniques. Essentially, the key difference from MAML lies in treating the prototypes of entity types in the support set as type labels for entities of the same type during training.
In the inner-loop update, the model first uses Equation (9) to predict the types for all true entity spans in the support set $S_{source}^{(i)}$ of each task. Then, the entity classification loss is computed based on $c^{N}_{k=1}(\mathcal{S})$ and Equation (10). Finally, gradient computation is performed using Equation (3) to optimize the parameters.
In the outer-loop update, the training process is similar to entity span detection. The model first uses Equation (9) to predict the types for all true entity spans in the query set $Q_{source}^{(i)}$ of each task. Then, based on $c^{N}_{k=1}(\mathcal{S})$ and Equation (10), the entity classification loss is computed. Subsequently, as shown in Equation (4), the losses across all tasks are accumulated. Finally, gradient computation is performed using Equation (5) to optimize the parameters.
Algorithm 2 Training Process of Entity Classification Model Based on MAML-ProtoNet Algorithm
@@ II-A2 Training
3:Randomly initialize $\theta$
4:while not done do
5:  Randomly select a batch of data $\mathcal{T}$, $\mathcal{T}\sim D_{source}$
6:  for all $\mathcal{T}$ do
7:   Select a task $(S_{source}^{(i)},Q_{source}^{(i)})$ from $\mathcal{T}$
8:   Obtain $c^{N}_{k=1}(\mathcal{S})$ by computing prototype vectors on $S_{source}^{(i)}$ using Equation (8)
9:   Based on $c^{N}_{k=1}(\mathcal{S})$ and the true entity spans in the support set $S_{source}^{(i)}$, compute $\nabla_{\theta}{L}_{\mathit{EC}}\left(S_{source}^{(i)};{\theta}\right)$ using Equations (9, 10)
10:   Perform gradient update: $\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{EC}}\left(S_{source}^{(i)};{\theta}\right)$
11:   Based on $c^{N}_{k=1}(\mathcal{S})$ and the true entity spans in the query set $Q_{source}^{(i)}$, compute $\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{EC}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ using Equations (9, 10) and save the results
12:  end for
13:  Outer-loop update: $\theta=\theta-\beta\sum\nolimits_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{EC}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$
14:end while
II-B Finetuning Process
Model finetuning aims to enable the model to adapt to new knowledge types during the domain transfer process. We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same pattern as training process.
II-C Inference Process
@@
(2) We input the query sentence $x^{\prime}$ into the finetuned best model for entity span detection (ESD) to get all entities. Specifically, we apply the Viterbi algorithm [17] for decoding, and derive the entities from the maximum sequence probability of Eq.2:
@@
TABLE I: F1 scores with standard deviations(%) on FewNERD. †denotes the results reported in [12] and [18]. ‡are the results we produce by LLM. * means the abbreviation of ‘contrastive learning’. The best results are in bold.
@@
TABLE II: The comparison of inference time between MsFNER and ChatGPT.
Models FewNERD-INTRA FewNERD-INTER FewAPTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot 1 shot 3 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way 4 way 6 way 4 way 6 way
MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266 0.5126 0.4981 0.7530 0.6055
ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409 2.3078 2.8610 4.9470 7.2926
(3) Firstly, we compute the representation $\hat{e}^{\prime}_{k}$ of each detected entity ${e}^{\prime}_{k}$ by Eq.6. Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ [19], and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.9.
@@
where the $\lambda$ is a hyper-parameter that makes a balance between KNN and prototype classifier.
III Experiments
III-A Experiments Setup